OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. (3 points) x2 − 25 = 0 and x2 = 2x + 15 Part 2: Describe what the solution(s) represent to the graph of each. (2 points) Part 3: How are the graphs alike? How are they different? (2 points)

OpenStudy (anonymous):

Part 1 A. x2 − 25 = 0 (x + 5) ( x - 5) = 0 x = 5, - 5 B. x2 = 2x + 15 x2 - 2x - 15 = 0 (x + 3) ( x - 5) = 0 x = 5, - 3 Part 2 A. the solution of x = 5, - 5 represents y values of the function to equal to zero and the only two points in the x-axis in which the function crosses. B. same reasoning, the solution of x = 5, - 3 represents y values of the function to equal to zero and the only two points in the x-axis in which the function crosses.

OpenStudy (anonymous):

hope tht helps :)

OpenStudy (anonymous):

omg thanks!! you did that yourself ?!